OpenStudy (anonymous):

True/False d/dx (sqrt(f(x)) = f '(x)/2sqrt(f(x))

OpenStudy (anonymous):

\[\frac{ d }{ dx } (\sqrt{f(x)}) = \frac{ f '(x) }{ 2\sqrt{f(x)} }\]

OpenStudy (anonymous):

When in doubt convert square roots to exponential notation, then the chain rule may become more clear: \[\frac{d}{dx}(f(x))^{\frac{1}{2}}=\frac{1}{2}(f(x))^{\frac{-1}{2}})*(f'(x))\]